The resultant of two forces in a plane is 400 N at 120 degrees. If one of the forces is 200 lbs at 20 degrees what is the other force?


1
Expert's answer
2021-08-23T12:39:34-0400

Given:

"R=400\:\rm N"

"\theta=120^{\circ}"

"F_1=200\:\rm lbs=890\: N"

"\alpha=20^{\circ}"




The resultant force

"\bf R=F_1+F_2"

Hence,

"\bf F_2=R-F_1"

"F_2=\sqrt{R^2+F_1^2-2RF_1\cos(120^{\circ}-20^{\circ})}"

"F_2=\sqrt{400^2+890^2-2*400*890\cos(100^{\circ})}\\
F_2=1037\:\rm N"